OpenStudy (neutrogena):

How do you expand (x - 11)^2?

OpenStudy (mikurout):

apply (a-b)^2 formula

OpenStudy (neutrogena):

What's the next step?

OpenStudy (518nad):

do u know how to expand a(b+c)

OpenStudy (neutrogena):

Do you distribute?

OpenStudy (518nad):

yes

OpenStudy (mikurout):

after applying that formula, u will get ur answer

OpenStudy (518nad):

can u expand a(b+c)

OpenStudy (neutrogena):

a(b+c) -> ab + ac...?

OpenStudy (518nad):

yes

OpenStudy (518nad):

so use that

OpenStudy (neutrogena):

Do I distribute with x?

OpenStudy (518nad):

(a+b)(a+b) = ?

OpenStudy (neutrogena):

I'm not sure, I skimmed through the lesson. :(

OpenStudy (neutrogena):

Is it aˆ2 + ab + ab + bˆ2 ?

OpenStudy (518nad):

:(

OpenStudy (518nad):

use what u already know

OpenStudy (518nad):

what do u do with y(a+b)

OpenStudy (neutrogena):

You distribute y with a and b.

OpenStudy (518nad):

write it out

OpenStudy (518nad):

we dont use english, we use math

OpenStudy (518nad):

y(a+b)=ya+yb y=(a+b) (a+b)(a+b)=(a+b)a+(a+b)b

OpenStudy (neutrogena):

Oh, alright!

OpenStudy (518nad):

(a+b)a+(a+b)b (aa+ba+ab+bb) a^2+2ab+b^2

OpenStudy (neutrogena):

Thank you :)

OpenStudy (518nad):

(x - 11)^2 (x-11)(x-11) (x - 11)x-(x - 11)11 xx-11x-(x11-11*11) x^2-11x - 11x +11^2 x^2-22x+11^2

OpenStudy (518nad):

:)

OpenStudy (neutrogena):

Thank you so much!

OpenStudy (518nad):

if you want to thank me, do this for me (X+3)(x+1)

OpenStudy (zzr0ck3r):

google "foil math expansion"

OpenStudy (neutrogena):

Sorry it's taking me a while, but would I have to do two separate expansions for that?

OpenStudy (518nad):

yeah so think of whatever in one bracket as just one variable

OpenStudy (518nad):

(x+3)(x+2) so lets make x+3 = a a(x+2) ax+2a now sub x+3=a (x+3)x+2(x+3)

OpenStudy (518nad):

expand the last line again